[figure: 9.7 cm]

THERMODYNAMICAL PROPERTIES OF THE SPIN-\frac{1}{2} HEISENBERG ANTIFERROMAGNET ON THE DECORATED SQUARE LATTICE

P. Tomczak^{a} and J. Richter^{b} aPhysics Department, Adam Mickiewicz University Umultowska 85, 61-614 Poznań, Poland bInstitut für Theoretische Physik, Otto-von-Guericke Universität Magdeburg P.O.B. 4120, 39016 Magdeburg, Germany

Two dimensional quantum Heisenberg spin systems with strong frustration which do not exhibit long range magnetic order in the ground state are currently being a subject of experimental and theoretical interest. It is well established that in such systems the excitation spectrum has a gap which can also be noticed in the temperature dependence of the specific heat. Recently it has been suggested^{1} (using 1/N expansion and Mean Field Approximation) that the Heisenberg antiferromagnet on the decorated (in a checkerboard way) square lattice (so-called squagome lattice) should have a disordered ground state and and two-peak specific heat. The aim of our work is to present a) the justification of the existence of the additional low-temperature specific-heat peak by means of quantum decimation scheme of Suzuki and Takano^{2} and b) some variational ground state functions for the Heisenberg antiferromagnet on the squagome lattice.

1Rahul Siddharthan, Antoine Georges, cond-mat/0103334. 2H. Takano, M. Suzuki, J. Stat. Phys. 26, 635 (1981).

[figure: 13.4 cm]

Subject category :

6. Theory of Magnetism

Presentation mode : poster

Corresponding author : P. Tomczak

Address for correspondence : Physics Department, Adam Mickiewicz University Umultowska 85, 61-614 Poznań, Poland

Email address : tomczak@man.poznan.pl